70 likes | 192 Vues
This document provides an update on the ongoing issues with the MFW, specifically the Phase Lock loss observed with the 53 MHz connected to the RFT at MI-10 on February 20, 2007. Despite attempts at resolution, the problem remains unresolved. A study plan is forthcoming to address this. Additionally, we report on the A0 Loss Monitor Testing in relation to the P-Bar Injection at 8 GeV and the new loss monitoring system's performance following a reset. Improvements in the LED signal for A0 loss monitoring are also highlighted.
E N D
FW Update 5/02/07 • MFW Latest problems • RFT Looses Phase Lock • Real 53 MHz connected to the RFT at MI-10 on Feb 20, 2007. • Apparently did not fix this problem • We will come up with a study plan • A0 Loss Monitor Testing
MFW Problem: P-Bar Injection 8Gev
MFW • P-Bar Injection 8Gev after reset